OpenStudy (anonymous):

Determine if x = 2 is a root for the function: f(x) = 2x3 - 22x2 + 76x - 80

OpenStudy (dan815):



OpenStudy (dan815):

if 2 is a root f(2) = 0 check if true

OpenStudy (anonymous):

how do i check tho

OpenStudy (dan815):

check f(2)

OpenStudy (dan815):

f(x) = 2x3 - 22x2 + 76x - 80 f(2) = put 2 where x is and simplify

OpenStudy (dan815):

f(x) = 2(x)3 - 22(x)2 + 76(x) - 80 f(2) = 2(2)3-22(2)2+76(2)-80

OpenStudy (dan815):

make sense?

OpenStudy (anonymous):

yeah thanks